OpenStudy (anonymous):

A rocket initially at rest accelerates at a rate of 99.0 meters/second2. Calculate the distance covered by the rocket if it attains a final velocity of 445 meters/second after 4.50 seconds. Please help im stuck and don't get it.

OpenStudy (akashdeepdeb):

S = ? a = 99m/s2 t = 4.5 s v = 445m/s Find u from equation (v=u+at) Then find S from (S = ut + 1/2 * a * t^2) Do you know the rocket-equation by any chance?